5.The kinetics of double strand formation for a DNA oligonucleotide containing a G.T base pair was measured by temperature-jump kinetics. The reaction is: 2CGTGAATTCGCG DUPLEX -1 The following data were obtained: Temperature (°C) 31.8 36.8 40.8 46.7 ki (M-ls-1 0.8x10-5 2.3x10-5 3.5x10-5 6.0x10-5 (A) Determine the equilibrium constant at each temperature from the forward and reverse rate constants. (B) Determine the free energy of the reaction at each temperature. (C) Using the forward reaction and the temperature dependence of k1, determine the pre-exponential factor and the energy of activation. Use a graphical method and assume that A and Ea are temperature 1.00 3.20 15.4 87.0 independent. (D) Repeat part C for the reverse reaction (i.e., the temperature dependence of
Answer #1

Griven Rencthon i k, Tempkev 31.8 36 .g 시0.8 46.7 0.3x Io 0-227 지。 o.069 xo ee enu 21.74 3o9.9 39227XI0 319 7 30.50 33 10 37.71 - Ea/RT 8.1 T Ea/RT a. d For reverse reactwn Temp (k) Зо 4-8 | О.ex10-5 309-8 2.3 XID 313-g | 3.5x® Data Table: o 0o328 73 o 3 20 o 00323 1o163 15 4 .0310260 731

Thus the equations from the ft are:

Forward reaction:

y = -13104 x + 31.432

slope = -Ea/R = -13104

Ea = 13104 * R = 13104*0.008314 = 108.95 kJ/mol

Intercept = lnA = 31.432

A = 4.47*10^13 M-1s-1

Reverse reaction

y = -30459x + 99.788

Ea = -30459*0.008314 = 253.24 kJ/mol

A = 2.17*10^43 M-1s-1